# Generative AI-enhanced Low-Altitude UAV-Mounted Stacked Intelligent Metasurfaces

### v1

@@ Abstract
Wireless communication systems face significant challenges in meeting the increasing demands for higher data rates and more reliable connectivity in complex environments. Stacked intelligent metasurfaces (SIMs) have emerged as a promising technology for realizing wave-domain signal processing, with mobile SIMs offering superior communication performance compared to their fixed counterparts. In this paper, we investigate a novel unmanned aerial vehicle (UAV)-mounted SIMs (UAV-SIMs) assisted communication system within the low-altitude economy (LAE) networks paradigm, where UAVs function as both base stations that cache SIM-processed data and mobile platforms that flexibly deploy SIMs to enhance uplink communications from ground users. To maximize network capacity, we formulate a UAV-SIM-based joint optimization problem (USBJOP) that comprehensively addresses three critical aspects: the association between UAV-SIMs and users, the three-dimensional positioning of UAV-SIMs, and the phase shifts across multiple SIM layers. Due to the inherent non-convexity and NP-hardness of USBJOP, we decompose it into three sub-optimization problems, i.e., association between UAV-SIMs and users optimization problem (AUUOP), UAV location optimization problem (ULOP), and UAV-SIM phase shifts optimization problem (USPSOP), and solve them using an alternating optimization strategy. Specifically, we transform AUUOP and ULOP into convex forms solvable by the CVX tool, while addressing USPSOP through a generative artificial intelligence (GAI)-based hybrid optimization algorithm. Simulations demonstrate that our proposed approach significantly outperforms benchmark schemes, achieving approximately 1.5 times higher network capacity compared to suboptimal alternatives. Furthermore, we reveal that SIMs effectively mitigate multi-user interference, with this effectiveness positively correlated with the number of SIM layers and meta-atoms per layer within specific ranges. Additionally, our proposed GAI method reduces the algorithm runtime by 10% while maintaining solution quality, thereby offering improved computational efficiency for practical deployments.
@@ I Introduction
Specifically, SIMs represent a significant advancement over traditional RIS technology by adopting a multi-layer architecture that mimics the structure of artificial neural networks [5]. This multi-layer configuration enables SIMs to perform sophisticated EM wave manipulations and computational tasks with significantly enhanced capabilities compared to single-layer structures [6]. A key advantage of SIMs is their ability to process signals at the speed of light through direct wave-domain computation, thereby eliminating the processing delays typically associated with conventional digital signal processing methods. Additionally, SIMs can achieve comparable beamforming performance while requiring fewer radio frequency chains, thus reducing the system dependency on high-precision analog-to-digital and digital-to-analog conversion components [4]. These characteristics make SIMs particularly attractive for applications requiring ultra-low latency and high computational efficiency in wireless networks.
@@
Despite their promising capabilities, SIM implementations face significant practical challenges when deployed with fixed base stations (BSs). Since SIM systems generally lack cache modules, they perform optimally when integrated with dedicated BSs [7]. However, fixed BS deployments inherently limit the potential of SIM technology by constraining coverage area, adaptability to changing network conditions, and ability to serve mobile users efficiently. These limitations are particularly pronounced in scenarios requiring rapid deployment, such as emergency response situations, temporary event coverage, or service provision in remote areas where fixed infrastructure is impractical or cost-prohibitive [8]. Furthermore, fixed deployments cannot easily adapt to dynamic user distributions and changing network traffic patterns, resulting in suboptimal resource utilization and reduced overall network performance [9].
@@
To overcome these limitations, mounting SIMs on mobile platforms within the LAE networks presents a compelling solution that could significantly enhance their versatility and performance. Among various mobile platform options, unmanned aerial vehicles (UAVs) have garnered particular interest due to their exceptional mobility, low deployment cost, and ability to establish favorable line-of-sight (LoS) communication channels with ground users [10]. As such, the superior LoS conditions facilitated by UAVs in the LAE networks are especially beneficial for SIM-based wave-domain calculations, as they reduce the complexity of channel estimation and improve the accuracy of signal processing operations. Moreover, UAVs possess a certain cache capacity, enabling them to store data processed through SIM-based computations, thus complementing the lack of storage capabilities of SIMs [11]. This synergistic integration creates the UAV-mounted SIMs (UAV-SIMs), which represent a promising technology for next-generation wireless networks.
@@
Thus, in this paper, we propose a comprehensive framework for UAV-SIM-assisted uplink communications that addresses these challenges through a novel alternating optimization method with generative artificial intelligence (GAI). Our key contributions are as follows:
@@
  • Performance Analysis: Simulation results demonstrate that UAV-SIMs utilizing the proposed AO strategy achieve approximately $1.5$ times the network capacity compared to suboptimal benchmarks. Moreover, the GAI method increases the overall speed of the AO algorithm by $10\%$ in complex scenarios, while ensuring the quality of the final solution. We further analyze the impact of key system parameters, including the number of SIM layers, the number of meta-atoms on each layer, and user density, and reveal that the SIM is highly effective at mitigating multi-user interference, and this effectiveness is positively correlated with the number of SIM layers and meta-atoms per layer within a specific range.
@@ II Related Works
In this section, we review the existing literature related to this paper on UAV-SIM-assisted uplink communications within LAE networks. Specifically, the literature review is organized into three main categories, including SIM architectures, UAV-assisted wireless communications, and optimization methods in SIM and UAV deployments.
@@ III-A System Overview
As depicted in Fig. 1, we consider an uplink communication scenario assisted by UAV-SIMs within the paradigm of LAE networks. Specifically, within a finite area, a group of UAV-SIMs =$\mathcal{M}=\{1,\ldots,M\}$ are deployed at a fixed altitude to serve a set of ground users $\mathcal{K}=\{1,\ldots,K\}$. Without loss of generality, we consider that $M<K$, which is common in practical applications where a limited number of UAVs are deployed to serve a larger user base. Given the physical constraints of actual deployment, such as limited payload, size, and power budget, each UAV-SIM is equipped with a single antenna. Meanwhile, to maximize spectrum utilization, particularly in scenarios with scarce spectrum, all UAV-SIMs operate on a shared frequency, which leads to mutual interference when multiple users communicate simultaneously. Additionally, the specific physical structure of the SIM is presented in Fig. 1, showing the stacked arrangement of multiple metasurface layers.
@@
Similar to [25], in order to ensure the user transmission speed across LAE networks, we consider that a UAV-SIM can only serve one user and a user can only be served by one UAV-SIM. Mathematically, we define a binary variable $S_{m,k}$, where $S_{m,k}=1$ means that there is an association between the $m$-th UAV-SIM and $k$-th user within LAE networks, i.e., $k$-th user can transmit data to the $m$-th UAV-SIM; otherwise, $S_{m,k}=0$, meaning that the $m$-th UAV-SIM does not provide service to the $k$-th user. The corresponding constraints can be written as follows:
@@ III-B SIM Transmission Model
Let =$\mathcal{L}=\{1,2,...,L\}$ represent the set of metasurface layers, and $\mathcal{N}=\{1,2,...,N\}$ represent the set of meta-atoms. Generally, we consider $\phi_{n}^{l}=\mathrm{e}^{j\theta_{n}^{l}}$ denotes the transmission coefficient imposed by the $n$-th meta-atom on the $l$-th metasurface layer, where $\theta_{n}^{l}$ denotes the corresponding phase shift, which satisfies $\theta_{n}^{l}\in[0,2\pi),\forall n\in\mathcal{N},\forall l\in\mathcal{L}$. Consequently, the diagonal phase shift matrix $\mathbf{\Phi}^{l}$ of the $l$-th metasurface layer is given as follows [36]:
@@
Next, we consider the transmission matrix between adjacent metasurfaces and from the transmitting antenna to the input metasurface of the SIM. Denote $\mathbf{W}^{l}\in\mathbb{C}^{N\times N},\forall l\in\mathcal{L}/\{1\}$ as the channel response from the $(l-1)$-th metasurface layer to the $l$-th metasurface layer, and $\mathbf{w}^{1\mathrm{H}}_{m}\in\mathbb{C}^{1\times N}$ is the channel response from the output metasurface layer of the SIM to the $m$-th UAV. According to the Rayleigh-Sommerfeld diffraction theory [37], the response from the $\tilde{n}$-th meta-atom on the $(l-1)$-th metasurface layer to the $n$-th meta-atom on the $l$-th metasurface layer can be written as follows [4]:
$w_{n,\tilde{n}}^{l}$ $=\frac{d_{x}d_{y}\cos\psi_{n,\tilde{n}}^{l}}{d_{n,\tilde{n}}^{l}}\left(\frac{1}{2\pi d_{n,\tilde{n}}^{l}}-j\frac{1}{\lambda}\right)\mathrm{e}^{j2\pi d_{n,\tilde{n}}^{l}/\lambda},$ (9)
$\quad\forall l\in\mathcal{L}\setminus\{1\},$
where $d_{n,\tilde{n}}^{l}$ is the transmission distance, $d_{x}d_{y}$ denotes the area of each meta-atom and $\psi_{n,\tilde{n}}^{l}$ is the angle between the propagation direction and the normal direction of the $(l-1)$-th metasurface layer. Similarly, the $n$-th entry $w^{1}_{n,m}$ of $\mathbf{w}^{1}_{m}$ can be obtained using Eq. (9).
@@ III-C UAV Energy Consumption Model
$\begin{split}P_{m}^{\text{prop}}&=P_{0}\left(1+\frac{3v_{m}^{2}}{U_{\text{tip}}^{2}}\right)+P_{i}\left(\sqrt{1+\frac{v_{m}^{4}}{4v_{0}^{4}}}-\frac{v_{m}^{2}}{2v_{0}^{2}}\right)^{1/2}\\ &+\frac{1}{2}d_{0}\rho sAv_{m}^{3},\end{split}$ (15)
@@ III-D Problem Formulation
[…]
@@ IV Proposed Solution
1 Let $\tau=0$ and initialize $\bm{\vartheta}$ and $\mathbf{W}$;
@@ IV-A1 AUUOP Formulation
[…]
@@ IV-B1 ULOP Formulation
[…]
@@ IV-B2 ULOP Solution
which is clearly a convex constraint.
@@ IV-C1 USPSOP Formulation
[…]
@@ IV-C2 USPSOP Solution
The core challenge of USPSOP fundamentally lies in the complex joint optimization of $L\times N$ atomic phase shifts. This optimization problem is inherently non-convex, making direct solutions extremely difficult. Given that UAV-SIMs and ground users in communication systems in LAE networks typically lack powerful computational capabilities, efficiently obtaining superior phase shift optimization results under limited resource constraints becomes key to solving USPSOP.
To effectively address the challenges posed by USPSOP, we propose an innovative hybrid generative phase shift optimization (HGPSO) algorithm. HGPSO can intelligently select and switch between two distinct sub-algorithms based on different communication scenarios and system configurations. This approach aims to achieve an optimal balance between optimization performance and computational efficiency. The two core sub-algorithms are the layer-by-layer iterative phase shift optimization (LBL-IPSO) algorithm [19] and a modified powerful generative artificial intelligence model, specifically a conditional variational autoencoder (CVAE) [15], namely CVAE-M. In the following sections, we will detail these two core sub-algorithms of HGPSO.

### v2

@@ Abstract
Wireless communication systems face challenges in meeting the demand for higher data rates and reliable connectivity in complex environments. Stacked intelligent metasurfaces (SIMs) have emerged as a promising technology for advanced wave-domain signal processing, where mobile SIMs can outperform fixed counterparts. In this paper, we propose a novel unmanned aerial vehicle (UAV)-mounted SIM (UAV-SIM) assisted communication system within low-altitude economy (LAE) networks, where UAVs act as both cache-enabled base stations and mobile SIM carriers to enhance uplink transmissions. To maximize network capacity, we formulate a UAV-SIM-based joint optimization problem (USBJOP) that integrates user association, UAV-SIM three-dimensional positioning, and multi-layer SIM phase shift design. Due to the non-convexity and NP-hardness of USBJOP, we decompose it into three subproblems, which are the association between UAV-SIMs and users optimization problem (AUUOP), the UAV location optimization problem (ULOP), and the UAV-SIM phase shifts optimization problem (USPSOP). Then, we solve them through an alternating optimization strategy. Specifically, AUUOP and ULOP are transformed into convex forms solvable via the CVX tool, while USPSOP is addressed by a generative artificial intelligence (GAI)-based hybrid optimization algorithm. Simulation results show that the proposed approach achieves approximately 1.5 times higher network capacity compared with suboptimal schemes, effectively mitigates multi-user interference with increasing SIM layers and meta-atoms, and reduces runtime by 10% while maintaining solution quality, thereby demonstrating its practicality for real-world deployments.
@@ I Introduction
Specifically, SIMs represent a significant advancement over traditional RIS technology by adopting a multi-layer architecture that mimics the structure of artificial neural networks [5]. Different from conventional single-layer RIS that merely reflects incident signals with adjustable phase shifts, the multi-layer SIM architecture employs multiple cascaded metasurface layers that exponentially increase the available degrees of freedom for electromagnetic wave manipulation. This multi-layer configuration enables SIMs to perform sophisticated EM wave manipulations and computational tasks at the speed of light through direct wave-domain computation, thereby eliminating the processing delays typically associated with conventional digital signal processing methods while achieving significantly enhanced capabilities compared to single-layer structures [6]. This capability is particularly crucial in mobile platform deployment scenarios where real-time adaptation to changing channel conditions is essential, as traditional single-layer RIS requires frequent beamforming optimization recalculations when platform positions change. Additionally, SIMs can achieve comparable beamforming performance while requiring fewer radio frequency chains, thus reducing the system dependency on high-precision analog-to-digital and digital-to-analog conversion components [4]. Furthermore, this enhanced control enables much more precise manipulation of wavefronts and effective multi-user interference suppression compared to the limited capabilities of single-layer designs. These characteristics make SIMs particularly attractive for applications requiring ultra-low latency and high computational efficiency in wireless networks.
@@
Despite their promising capabilities, SIM implementations face significant practical challenges when deployed with fixed base stations (BSs). Since SIM systems generally lack cache modules, they perform optimally when integrated with dedicated BSs [7, 8]. However, fixed BS deployments inherently limit the potential of SIM technology by constraining coverage area, adaptability to changing network conditions, and ability to serve mobile users efficiently. These limitations are particularly pronounced in scenarios requiring rapid deployment, such as emergency response situations, temporary event coverage, or service provision in remote areas where fixed infrastructure is impractical or cost-prohibitive [9]. Furthermore, fixed deployments cannot easily adapt to dynamic user distributions and changing network traffic patterns, resulting in suboptimal resource utilization and reduced overall network performance [10].
@@
To overcome these limitations, mounting SIMs on mobile platforms within the LAE networks presents a compelling solution that could significantly enhance their versatility and performance. Among various mobile platform options, unmanned aerial vehicles (UAVs) have garnered particular interest due to their exceptional mobility, low deployment cost, and ability to establish favorable line-of-sight (LoS) communication channels with ground users [11]. In particular, several application domains can significantly benefit from UAV-assisted mobile communication platforms. For example, in emergency response operations such as earthquakes or floods where terrestrial communication infrastructure is often severely damaged, UAV-based systems can be rapidly deployed to reconfigure coverage and maintain connectivity for geographically dispersed survivors [12]. In the agricultural sector, UAV-assisted platforms can flexibly patrol farmland and dynamically extend coverage to temporary blind spots or high-demand areas, thereby supporting high-resolution sensing and imagery transmission [13]. Similarly, in urban environments characterized by dense user distributions and severe signal blockage from high-rise buildings, UAV-based mobile platforms can adaptively adjust both positions and communication parameters to enhance traffic management and public safety services [14]. As such, the superior LoS conditions facilitated by UAVs in the LAE networks are especially beneficial for SIM-based wave-domain calculations, as they reduce the complexity of channel estimation and improve the accuracy of signal processing operations. Moreover, UAVs possess a certain cache capacity, enabling them to store data processed through SIM-based computations, thus complementing the lack of storage capabilities of SIMs [15]. This synergistic integration creates the UAV-mounted SIMs (UAV-SIMs), which represent a promising technology for next-generation wireless networks.
@@
Thus, in this paper, we propose a comprehensive framework for UAV-SIM-assisted uplink communications that addresses these challenges through a novel alternating optimization method with generative artificial intelligence (GAI) [18]. Our key contributions are as follows:
@@
  • Performance Analysis: Simulation results demonstrate that UAV-SIMs utilizing the proposed AO strategy achieve approximately $1.5$ times the network capacity compared to suboptimal benchmarks. Moreover, the GAI method increases the overall speed of the AO algorithm by $10\%$ in complex scenarios, while ensuring the quality of the final solution. We further analyze the effectiveness of our designed loss function with the capacity-oriented term, which maintains consistent performance advantages ranging from $20\%$ to $400\%$ across all configurations. In addition, we investigate the impact of key system parameters, including the number of SIM layers, the number of meta-atoms on each layer, and user density, revealing that the SIM is highly effective at mitigating multi-user interference, and this effectiveness is positively correlated with the number of SIM layers and meta-atoms per layer within a specific range. Finally, we expand the simulation scenarios and find that the system continues to exhibit these favorable characteristics.
@@ II Related Works
In this section, we review the existing literature related to this paper on UAV-SIM-assisted uplink communications within LAE networks. Specifically, the literature review is organized into three main categories, including SIM architectures, UAV-assisted wireless communications, and optimization methods in SIM and UAV deployments. To further highlight the differences between this paper and previous works, we summarize a comparison table as shown in Table I.
@@ II-C Optimization Methods in SIM and UAV Deployments
TABLE I: Comparison between related works and our contribution
Reference
UAV
SIM/RIS
Optimization
Method
Communication
Metrics
User
Association
3D
Positioning
SIM/RIS
Phase Shifts
[21] $\times$ $\checkmark$ Convex Optimization Network Capacity $\checkmark$ $\times$ $\checkmark$
[22] $\times$ $\checkmark$ Convex Optimization Transmit Power $\checkmark$ $\times$ $\checkmark$
[4] $\times$ $\checkmark$ Traditional Error Minimization $\times$ $\times$ $\checkmark$
[25] $\times$ $\checkmark$ Traditional DOA Estimation $\times$ $\times$ $\checkmark$
[26] $\checkmark$ $\times$ Convex Optimization Sum Rate $\checkmark$ $\checkmark$ $\times$
[27] $\checkmark$ $\times$ Learning-based Energy Efficiency $\checkmark$ $\checkmark$ $\times$
[16] $\checkmark$ $\times$ Convex Optimization Min Throughput $\checkmark$ $\checkmark$ $\times$
[32] $\checkmark$ $\times$ Heuristic Task Completion $\checkmark$ $\times$ $\times$
[33] $\times$ $\checkmark$ Traditional Spectral Efficiency $\checkmark$ $\times$ $\checkmark$
[36] $\checkmark$ $\times$ Traditional User Throughput $\checkmark$ $\checkmark$ $\times$
[38] $\checkmark$ $\checkmark$ Traditional Coverage Probability $\checkmark$ $\times$ $\times$
[39] $\checkmark$ $\checkmark$ Traditional System Capacity $\times$ $\times$ $\times$
This Work $\checkmark$ $\checkmark$ Convex Optimization Sum Rate $\checkmark$ $\checkmark$ $\checkmark$
@@ III-A System Overview
As depicted in Fig. 1, we consider an uplink communication scenario assisted by UAV-SIMs within the paradigm of LAE networks. Specifically, within a finite area, a group of UAV-SIMs $\mathcal{M}=\{1,\ldots,M\}$ are deployed at a fixed altitude to serve a set of ground users $\mathcal{K}=\{1,\ldots,K\}$. Without loss of generality, we consider that $M<K$, which is common in practical applications where a limited number of UAVs are deployed to serve a larger user base. Given the physical constraints of actual deployment, such as limited payload, size, and power budget, each UAV-SIM is equipped with a single antenna. Meanwhile, to maximize spectrum utilization, particularly in scenarios with scarce spectrum, all UAV-SIMs operate on a shared frequency, which leads to mutual interference when multiple users communicate simultaneously. Additionally, the specific physical structure of the SIM is presented in Fig. 1, showing the stacked arrangement of multiple metasurface layers.
@@
Similar to [16], we consider that a UAV-SIM can serve one user and a user can only be served by one UAV-SIM. Note that this one-to-one association model is adopted to establish a foundational framework for the UAV-SIM assisted communication system, which allows us to clearly delineate the core optimization challenges without the added complexity of multi-user resource allocation mechanisms and enables precise analysis of the impact of various system parameters on network capacity. Furthermore, this model can be readily extended to accommodate multi-user scenarios through time-division multiple access [16] schemes where a single UAV-SIM serves multiple users in different time slots by reconfiguring its phase shifts accordingly. As such, for the considered model, mathematically we define a binary variable $S_{m,k}$, where $S_{m,k}=1$ means that there is an association between the $m$-th UAV-SIM and $k$-th user within LAE networks, i.e., $k$-th user can transmit data to the $m$-th UAV-SIM; otherwise, $S_{m,k}=0$, meaning that the $m$-th UAV-SIM does not provide service to the $k$-th user. The corresponding constraints can be written as follows:
@@ III-B SIM Transmission Model
Let $\mathcal{L}=\{1,2,...,L\}$ represent the set of metasurface layers, and $\mathcal{N}=\{1,2,...,N\}$ represent the set of meta-atoms. Generally, we consider $\phi_{n}^{l}=\mathrm{e}^{j\theta_{n}^{l}}$ denotes the transmission coefficient imposed by the $n$-th meta-atom on the $l$-th metasurface layer, where $\theta_{n}^{l}$ denotes the corresponding phase shift, which satisfies $\theta_{n}^{l}\in[0,2\pi),\forall n\in\mathcal{N},\forall l\in\mathcal{L}$. Consequently, the diagonal phase shift matrix $\mathbf{\Phi}^{l}$ of the $l$-th metasurface layer is given as follows [40]:
@@
Next, we consider the transmission matrix between adjacent metasurfaces and from the transmitting antenna to the input metasurface of the SIM. Denote $\mathbf{W}^{l}\in\mathbb{C}^{N\times N},\forall l\in\mathcal{L}/\{1\}$ as the channel response from the (l1)(l-1)-th metasurface layer to the $l$-th metasurface layer, and $\mathbf{w}^{1\mathrm{H}}_{m}\in\mathbb{C}^{1\times N}$ is the channel response from the output metasurface layer of the SIM to the $m$-th UAV. According to the Rayleigh-Sommerfeld diffraction theory [41], the response from the $\tilde{n}-$th meta-atom on the (l1)(l-1)-th metasurface layer to the $n$-th meta-atom on the $l$-th metasurface layer can be written as follows [4]:
$w_{n,\tilde{n}}^{l}=\frac{d_{x}d_{y}\cos\psi_{n,\tilde{n}}^{l}}{d_{n,\tilde{n}}^{l}}\left(\frac{1}{2\pi d_{n,\tilde{n}}^{l}}-j\frac{1}{\lambda}\right)\mathrm{e}^{j2\pi d_{n,\tilde{n}}^{l}/\lambda},\quad\forall l\in\mathcal{L}\setminus\{1\},$ (9)
where $d_{n,\tilde{n}}^{l}$ is the transmission distance, $d_{x}d_{y}$ denotes the area of each meta-atom and $\psi_{n,\tilde{n}}^{l}$ is the angle between the propagation direction and the normal direction of the (l1)(l-1)-th metasurface layer. Similarly, the $n$-th entry $w^{1}_{n,m}$ of $\mathbf{w}^{1}_{m}$ can be obtained using Eq. (9).
@@ III-C UAV Energy Consumption Model
$\begin{split}P_{m}^{\text{prop}}=P_{0}\left(1+\frac{3v_{m}^{2}}{U_{\text{tip}}^{2}}\right)+P_{i}\left(\sqrt{1+\frac{v_{m}^{4}}{4v_{0}^{4}}}-\frac{v_{m}^{2}}{2v_{0}^{2}}\right)^{1/2}+\frac{1}{2}d_{0}\rho sAv_{m}^{3},\end{split}$ (15)
@@ III-D Problem Formulation
$\text{(USBJOP) }:\underset{\mathrm{\mathbf{S},\mathbf{W},\bm{\vartheta}}}{\max}\quad$ $R$ (20a)
@@ IV Proposed Solution
Figure 2: The structure of the AO strategy framework in LAE networks.
1Let $\tau=0$ and initialize $\bm{\vartheta}$ and $\mathbf{W}$;
@@ IV-A1 AUUOP Formulation
$\text{(AUUOP)}:\underset{\mathrm{\mathbf{S}}}{\max}\quad$ $R$ (21a)
@@ IV-B1 ULOP Formulation
$\text{(ULOP)}:\underset{\mathrm{\mathbf{W}}}{\max}\quad$ $R$ (23a)
@@ IV-B2 ULOP Solution
which is clearly a convex constraint. This energy constraint $\mathcal{C}_{10}$ is enforced as a hard constraint within each iteration of our alternating optimization procedure, which ensures that every UAV position update remains within the energy budget and makes the solutions immediately applicable to real-world deployments.
@@ IV-C1 USPSOP Formulation
$\text{ (USPSOP) }:\underset{\mathrm{\bm{\vartheta}}}{\max}\quad$ $R$ (34a)
@@ IV-C2 USPSOP Solution
The core challenge of USPSOP fundamentally lies in the complex joint optimization of $L\times N$ atomic phase shifts, which presents significantly greater complexity than conventional single-layer RIS optimization. Specifically, the operational mechanism of SIM itself introduces substantial mathematical complexity through its cascaded modulation architecture, where the output of each layer serves as the input to the subsequent layer, creating strong interdependencies among phase variables that prevent the direct application of traditional RIS optimization methods. This architectural difference results in exponential growth in computational complexity, while single-layer RIS optimization involves only $N$ independent phase variables, multi-layer SIM requires the simultaneous optimization of $L\times N$ highly coupled variables. Moreover, the resulting optimization problem is inherently non-convex, making direct analytical solutions intractable. These theoretical challenges are further compounded by practical constraints, as UAV-SIMs and ground users in LAE communication networks typically operate under severe computational limitations. Consequently, developing computationally efficient algorithms that can achieve superior phase shift optimization performance within these resource constraints becomes the key to solving USPSOP effectively.
To effectively address the challenges posed by USPSOP, we propose an innovative hybrid generative phase shift optimization (HGPSO) algorithm. HGPSO can intelligently select and switch between two distinct sub-algorithms based on different communication scenarios and system configurations. This approach aims to achieve an optimal balance between optimization performance and computational efficiency. The two core sub-algorithms are the layer-by-layer iterative phase shift optimization (LBL-IPSO) algorithm [24] and a modified powerful GAI model [47], specifically a conditional variational autoencoder (CVAE) [20], namely CVAE-M. By combining these two methods, HGPSO enhances exploration capability and accelerates convergence while maintaining solution quality, making it more suitable for addressing phase configuration problems of SIM in dynamic environments. In the following sections, we will detail these two core sub-algorithms of HGPSO.